the amount of an investment of P dollars for t years at simple interest rate r is given by A=P+Prt.
A) rewrite this formula by factoring out the GCF on the right-hand side.
B) find A if $830 is invested for 3 years at a simple interest rate of 15%.

Answer by math_il(6)

A) GCF is P
so A = P(1+rt)
B)P = 830 t=3y r=15%
substitute in the formula
A= 830( 1+ (15/100*3) )
simplifying
A = 830( 1 + 0.45) = 830*1.45 = 1203.5 $